# Angular-momentum pairs in spherical systems: applications to the Galactic centre

### v1

@@ 1 Introduction
The general $N$-body problem of self-gravitating systems does not have a closed-form solution; however, when the system has certain symmetries or a hierarchical structure with well-separated time-scales, the long-time dynamics become tractable and amenable to detailed study (see, e.g. Heggie & Hut, 2003, for a review of the $N$-body problem). In these systems, the underlying assumption is that the short-term orbital motions can be averaged out – this is the basis of the orbit-averaging method in Hamiltonian perturbation theory (e.g. Arnold et al., 2006). By averaging over the rapid orbital period, one obtains a simplified “secular” Hamiltonian that governs the long-term evolution of the system’s orbital elements, such as the orbital energy and angular momentum.
[…]
@@
This time-scale hierarchy allows one to employ a double orbit-averaging procedure: first, one averages over the rapid orbital period, replacing the time-varying orbit with a stationary eccentric wire, and then one also averages over the longer apsidal-precession period induced by the smooth spherical stellar potential, replacing the planar-precessing orbit with rings or two-dimensional annuli, whose shapes are preserved (because semi-major axes and eccentricities are conserved over the relevant time-scales). The gravitational torques between these rings drive a slow change of their orientations; the relevant dynamical variables are thus the directions of orbital angular momenta, $\hat{\bm{L}}$, which are set by the inclination and the argument of the ascending node. This formalism is referred to as “vector resonant relaxation” (VRR) (Rauch & Tremaine, 1996; Kocsis & Tremaine, 2011, 2015; Bar-Or & Fouvry, 2018; Fouvry et al., 2022; Panamarev & Kocsis, 2022). The VRR equations of motion can be solved exactly for the case of two rings; this results in a precession of the angular-momentum vectors of the two bodies about their total angular momentum, with a frequency which depends on the relative orbital configuration of the bodies (Kocsis & Tremaine, 2015).
@@
This problem is directly relevant to galactic nuclei: in the Milky Way, which hosts a super-massive black hole of mass $M=4\times 10^{6}\,{\rm M}_{\odot}$ (Ghez et al., 2000; Genzel et al., 2010), the nuclear star cluster (NSC) is dominated by old stars but also contains a population of young stars, some of which reside in a disc-like configuration (e.g. Levin & Beloborodov, 2003; Paumard et al., 2006; von Fellenberg et al., 2022; Jia et al., 2023). Initial evidence for this disc arose from the clustering of stellar angular-momentum vector directions $\hat{\bm{L}}$ (Levin & Beloborodov, 2003); however, recent analyses reveal a more complex distribution featuring multiple sub-clusters of these vectors (Bartko et al., 2009; von Fellenberg et al., 2022), on the sphere of $\hat{\bm{L}}$. Although certain sub-structures remain under debate (Jia et al., 2023), there is broad agreement on at least two distinct angular momentum clusters, one of which is retrograde relative to the other (von Fellenberg et al., 2022; Jia et al., 2023). A natural formation scenario invokes the fragmentation of a past gaseous accretion disc (Levin & Beloborodov, 2003), but the presence of retrograde stars suggests additional processes – such as perturbations by an external massive perturber or torques from a rotating NSC (cf. Levin, 2024). The young stellar disc is estimated to be $\sim 6$ Myr old (Genzel et al., 2010; Habibi et al., 2017), comparable to the VRR time-scales at these radii (Kocsis & Tremaine, 2011). Consequently, resonant relaxation has been widely invoked to explain the observed distribution of stellar angular-momentum vectors under the influence of the NSC (Kocsis & Tremaine, 2011, 2015; Roupas et al., 2017; Giral Martínez et al., 2020; Fouvry et al., 2019, 2022; Magnan et al., 2022; Panamarev & Kocsis, 2022; Wang & Kocsis, 2023; Flores & Fouvry, 2025), as well as their modification by an external perturber (Szölgyén et al., 2021; Ginat et al., 2023; Fouvry et al., 2023; Levin, 2024). Although VRR is a promising mechanism for producing the observed misalignments, other relaxation processes – such as scalar resonant relaxation (Bar-Or & Fouvry, 2018) and two-body (non-resonant) relaxation – may also contribute to the dynamics (see e.g. Panamarev & Kocsis 2022, who showed that two-body relaxation dominates for galactic nuclei with a light SMBH). The most accurate way to treat the simultaneous action of all of these processes is with high-resolution direct $N$-body simulations, which integrate the exact gravitational forces without relying on the orbit-averaged VRR approximation; however, direct $N$-body simulations of the full NSC remain computationally prohibitive – state-of-the-art models of the entire nuclear cluster require months to complete, even on modern GPUs (Panamarev et al., 2019; Panamarev & Kocsis, 2022).
@@ 2 Secular VRR Hamiltonian and Asymptotic Limits
We begin by considering a system of $N$ objects moving on apsidally precessing Keplerian orbits around a central SMBH, with fixed semi-major axes $a_{i}$ and eccentricities $e_{i}$. The central mass is denoted $M$. We assume that $\sum_{i}m_{i}\ll M$, so the gravitational forces from the stars on each other can be treated as small perturbations on otherwise Keplerian orbits. General relativistic corrections and external torques (e.g. from a massive perturber or frame-dragging) are neglected or assumed sub-dominant on the time-scales of interest, unless stated otherwise. Within the VRR framework, over time-scales long compared to an orbital period but short compared to changes in the scalar angular momenta $L_{i}$, the orbits can, therefore, be treated as apsidally precessing annuli whose mutual gravitational interaction is suitably averaged (see e.g. Kocsis & Tremaine, 2015). To build an intuitive understanding of these interactions, we begin by analysing the mutual torques between two circular rings; we then describe the Hamiltonian of the system, which provides a compact and general framework for describing their long-term secular evolution.
@@ 2.1 Torques between rings
where $\eta_{1}=m_{1}/2\pi R_{1}$ and $\eta_{2}=m_{2}/2\pi R_{2}$ are the linear mass densities of ring 1 and ring 2, respectively. Here, $m_{1}$ and $m_{2}$ are the total masses of rings 1 and 2, respectively. The integrals are taken over $\phi_{1}$ and $\phi_{2}$, which represent the azimuthal angles in the coordinate systems corresponding to ring 1 and ring 2. This equation reveals that co-planar rings ($I=0$) exert no gravitational torque on each other, as $\bm{r}_{1}\times\bm{r}_{2}=0$. Similarly, for polar orbits, when $I=\pi/2$, the integral cancels out due to symmetrical contributions from opposite points on ring 2; the system’s symmetry also leads to $\left(\tau_{12}\right)_{z}=0$. Thus, if we now take the rings to represent a star and an IMBH, with $m_{2}\gg m_{1}$, the $x$-component of the torque induces nodal precession of the star’s angular momentum vector around that of the IMBH.
While Eq. (1) offers an intuitive understanding of the gravitational interactions between two rings, extending this analysis to the $N$-ring system as a whole, and generalising for elliptical orbits, are more conveniently achieved through the Hamiltonian formalism. The gravitational interaction-energy of the ring $i$ with all other rings (Kocsis & Tremaine, 2015) can be written as:
@@
$\mathcal{J}_{ij\ell}=\frac{Gm_{i}m_{j}}{a_{\rm out}}P_{\ell}(0)^{2}\,s_{ij\ell}\,\alpha_{ij}^{\ell}\,,$ (3)
where $\alpha_{ij}\equiv a_{\mathrm{in}}/a_{\mathrm{out}}$, $a$ is the semi-major axis with indices ‘${\mathrm{out}}$’ and ‘${\mathrm{in}}$’ denoting the larger and smaller semi-major axes among $i$ and $j$, respectively; the coefficient $s_{ij\ell}$ is:
@@ 2.2 Equation of motion and asymptotic limits
From the Hamiltonian (2), one can get the equation of motion (Kocsis & Tremaine, 2015):
@@
$\Omega_{i,j}$ $=2\pi\omega_{i}\frac{m_{j}}{M}\left(\frac{r_{i}}{r_{\mathrm{out}}}\right)\left(\frac{r_{\mathrm{in}}}{r_{\mathrm{out}}}\right)^{2}P^{2}_{2}(0)P^{\prime}_{2}(\cos{\theta})$
@@
$M$ is the mass of the SMBH, $\Omega_{i,j}\neq\Omega_{ji}$, but $\Omega_{i,j}\times L_{i}=\Omega_{ji}\times L_{j}$, and $\omega_{i}=(GM/a_{i}^{3})^{1/2}$ is the orbital frequency around the SMBH and $\theta$ is the angle between the angular-momentum vectors given by $\cos\theta=\hat{\bm{L}}_{i}\cdot\hat{\bm{L}}_{j}$. This expression is a good approximation when $r_{\rm in}/r_{\rm out}\lesssim 0.3$, as we show in appendix C. One can use an asymptotic expansion (as $\ell\to\infty$) to derive an accurate approximation for arbitrary $r_{\rm in}/r_{\rm out}$; we do so explicitly in Appendix C.
@@
On the other hand, for orbits with close separations which include overlapping or marginally overlapping orbits, large $\ell$ harmonics dominate, because the coefficient $\mathcal{J}_{ij\ell}$ decays slowly with $\ell$, and the torques can be approximated by asymptotic expressions for $\mathcal{J}_{ij\ell}$ at $\ell\to\infty$. As we show in Appendices A and B (cf. Kocsis & Tremaine, 2015) an asymptotic expression valid for both overlapping ($z=1$) and non-overlapping orbits ($z<1$) is
@@
=$=\left\{\begin{array}[]{ll}\dfrac{m_{j}a_{i}}{\pi^{2}Mr_{\rm p,out}}\dfrac{[(1+e_{\rm in})(1-e_{\rm out})]^{3/2}}{(e_{i}e_{j})^{1/2}(1-e_{i}^{2})^{1/2}}&\text{if non-overlapping}\\[8.61108pt] \dfrac{4}{\pi^{3}}\dfrac{m_{j}}{M}\dfrac{I_{2}}{a_{j}(1-e_{i}^{2})^{1/2}}&\text{if overlapping}\end{array}\right.$ (14)
Here, $I_{2}\equiv I^{(2)}(r_{\rm p,in},r_{\rm p,out},r_{\rm a,in},r_{\rm a,out})$ is an analytic function expressed with elliptic integrals in a closed form, given by Eq. (68) of Kocsis & Tremaine (2015) which simplifies for the case of a circular ring embedded within an elliptical one, as we show in Appendix C.2. In Appendix C.1 we show how to construct equations of motion valid for arbitrary values of radial separation of the orbits.
In § 3 below, we will use these asymptotic expansions, together with exact $-\ell$ terms, to derive and compare analytical torque-estimates in systems influenced by an additional massive perturber.
@@ 3 Relative torques
$\left|\mathbf{\Omega}_{i,j}\times\bm{L}_{i}\right|\leq\left|\left(\mathbf{\Omega}_{i,\bullet}\times\bm{L}_{i}-\mathbf{\Omega}_{j,\bullet}\times\bm{L}_{j}\right)\cdot\Delta\hat{\bm{L}}\right|$ (15)
@@
$\left|\mathbf{\Omega}_{i,j}\times\bm{L}_{i}\right|\leq$ $\left|\left[\mathbf{\Omega}_{\bullet}\left(\theta_{\bullet}+\frac{\theta_{b}}{2},a+\frac{\Delta a}{2},e+\frac{\Delta e}{2}\right)\times\bm{L}_{i}\right.\right.$
$-\left.\left.\mathbf{\Omega}_{\bullet}\left(\theta_{\bullet}-\frac{\theta_{b}}{2},a-\frac{\Delta a}{2},e-\frac{\Delta e}{2}\right)\times\bm{L}_{j}\right]\cdot\Delta\hat{\bm{L}}\right|.$ (16)
@@
-- $\frac{1}{2}\left(\frac{\partial\mathbf{\Omega}_{\bullet}}{\partial\theta}\theta_{b}+\frac{\partial\mathbf{\Omega}_{\bullet}}{\partial a}\Delta a+\frac{\partial\mathbf{\Omega}_{\bullet}}{\partial e}\Delta e\right)\times(\bm{L}_{i}+\bm{L}_{j})\Bigg{]}\cdot\Delta\hat{\bm{L}}\Bigg{|}.$
@@
×$\times|\bm{L}_{i}+\bm{L}_{j}|\sin\theta_{\bullet},$ (18)
where $\kappa$ is a factor of order unity which accounts for the change in the orientation of the relative angular momentum vector of the pair with respect to the perturber. The geometry of the system gives constrains on the maximum relative inclination of the pair – when the system remains coupled, the relative angle is constant; but when it is decoupled, the relative angle varies between $\theta_{b}=\max|\theta_{1}\pm\theta_{2}|$ where “max” stands for the configuration where the initial polar angles $\theta_{1,2}$ are aligned according to a meridian as in Figure 1(a).
@@
We now restrict our attention to a special – but analytically tractable – configuration, in which the two bodies share the same semi-major axis and eccentricity, viz. $a_{i}=a_{j}=a$ and $e_{i}=e_{j}=e$. In such a synchronous pair, disruption can arise only from changes in their mutual inclination. The stability criterion therefore reduces to
$\Omega_{i,j}\!\left(\theta_{b},1\right)m_{i}\leq\;\frac{1}{2}\kappa\,\sin\theta_{\bullet}\,\frac{\partial\Omega_{\bullet}\!\left(\theta_{\bullet},a/a_{\bullet}\right)}{\partial\theta}(m_{i}+m_{j}),$ (19)

### v2

@@ 1 Introduction
The general $N$-body problem of self-gravitating systems does not have a closed-form solution; however, when the system has certain symmetries or a hierarchical structure with well-separated time-scales, the long-time dynamics become tractable and amenable to detailed study (see, e.g. Heggie and Hut, 2003, for a review of the $N$-body problem). In these systems, the underlying assumption is that the short-term orbital motions can be averaged out – this is the basis of the orbit-averaging method in Hamiltonian perturbation theory (e.g. Arnold et al., 2006). By averaging over the rapid orbital period, one obtains a simplified “secular” Hamiltonian that governs the long-term evolution of the system’s orbital elements, such as the orbital energy and angular momentum.
Several astrophysical systems allow for such a simplification, including planetary systems, galactic discs, and stellar clusters, such as globular clusters and nuclear star clusters. In the case of star clusters, they are approximately spherically symmetric, implying that to first order (i.e. the monopole in a multipole expansion) the stars in them move on planar orbits with conserved orbital energy and angular momentum in the smooth mean-field gravitational potential of the system. Over longer time-scales, the orbital energy and angular-momentum vectors slowly diffuse due to the fluctuating component of the potential or close two-body encounters (Binney and Tremaine, 2008). A system of $N$ bodies bound to a massive central body, including planets around a star in a planetary system or stars around a super-massive black hole (SMBH) in a nuclear star cluster, evolves similarly since the Keplerian potential is spherically symmetric. In nuclear star clusters, all objects perform Keplerian eccentric orbits around the central SMBH with nearly fixed period ($\sim 10$$10^{4}\,\mathrm{yr}$ in the case of the Milky Way, e.g. Kocsis and Tremaine 2011), semi-major axis $a$, eccentricity $e$, orbital inclination $i$, and argument of ascending node $\Omega$, corresponding to fixed orbital energy $E$ and angular-momentum vector $\bm{L}$. These eccentric orbits additionally execute rapid in-plane apsidal precession, due to the mean-field potential of the stellar distribution on time-scales of $\sim 10^{4}\,\mathrm{yr}$, longer than the orbital period, covering two-dimensional rings (or annuli) between their respective periapsis and apoapsis. On even longer time-scales, the direction of the orbital planes or angular-momentum vectors reorient ($\sim 10^{6}\,\mathrm{yr}$) (Rauch and Tremaine, 1996; Kocsis and Tremaine, 2015). The eccentricities and semi-major axes change on average even more slowly ($\sim 10^{8}\,\mathrm{yr}$ and $\sim 10^{9}\,\mathrm{yr}$, respectively).
@@
This time-scale hierarchy allows one to employ a double orbit-averaging procedure: first, one averages over the rapid orbital period, replacing the time-varying orbit with a stationary eccentric wire, and then one also averages over the longer apsidal-precession period induced by the smooth spherical stellar potential, replacing the planar-precessing orbit with rings or two-dimensional annuli, whose shapes are preserved (because semi-major axes and eccentricities are conserved over the relevant time-scales). The gravitational torques between these rings drive a slow change of their orientations; the relevant dynamical variables are thus the directions of orbital angular momenta, $\hat{\bm{L}}$, which are set by the inclination and the argument of the ascending node. This formalism is referred to as “vector resonant relaxation” (VRR) (Rauch and Tremaine, 1996; Kocsis and Tremaine, 2011, 2015; Bar-Or and Fouvry, 2018; Fouvry et al., 2022; Panamarev and Kocsis, 2022). The VRR equations of motion can be solved exactly for the case of two rings; this results in a precession of the angular-momentum vectors of the two bodies about their total angular momentum, with a frequency which depends on the relative orbital configuration of the bodies (Kocsis and Tremaine, 2015).
@@
This problem is directly relevant to galactic nuclei: in the Milky Way, which hosts a super-massive black hole of mass $M=4.3\times 10^{6}\,{\rm M}_{\odot}$ (Ghez et al., 2000; Genzel et al., 2010; Gravity Collaboration et al., 2024), the nuclear star cluster (NSC) is dominated by old stars but also contains a population of young stars, some of which reside in a disc-like configuration (e.g. Levin and Beloborodov, 2003; Paumard et al., 2006; von Fellenberg et al., 2022; Jia et al., 2023). Initial evidence for this disc arose from the clustering of stellar angular-momentum vector directions $\hat{\bm{L}}$ (Levin and Beloborodov, 2003); however, recent analyses reveal a more complex distribution featuring multiple sub-clusters of these vectors (Bartko et al., 2009; von Fellenberg et al., 2022), on the sphere of $\hat{\bm{L}}$. Although certain sub-structures remain under debate (Jia et al., 2023), there is broad agreement on at least two distinct angular momentum clusters, one of which is retrograde relative to the other (von Fellenberg et al., 2022; Jia et al., 2023). A natural formation scenario invokes the fragmentation of a past gaseous accretion disc (Levin and Beloborodov, 2003), but the presence of retrograde stars suggests additional processes – such as perturbations by an external massive perturber or torques from a rotating NSC (cf. Levin, 2024). The young stellar disc is estimated to be $\sim 6$ Myr old (Genzel et al., 2010; Habibi et al., 2017), comparable to the VRR time-scales at these radii (Kocsis and Tremaine, 2011). Consequently, resonant relaxation has been widely invoked to explain the observed distribution of stellar angular-momentum vectors under the influence of the NSC (Kocsis and Tremaine, 2011, 2015; Roupas et al., 2017; Giral Martínez et al., 2020; Fouvry et al., 2019, 2022; Magnan et al., 2022; Panamarev and Kocsis, 2022; Wang and Kocsis, 2023; Flores and Fouvry, 2025), as well as their modification by an external perturber (Szölgyén et al., 2021; Ginat et al., 2023; Fouvry et al., 2023; Levin, 2024). Although VRR is a promising mechanism for producing the observed misalignments, other relaxation processes – such as scalar resonant relaxation (Bar-Or and Fouvry, 2018) and two-body (non-resonant) relaxation – may also contribute to the dynamics (see e.g. Panamarev and Kocsis 2022, who showed that two-body relaxation dominates for galactic nuclei with a light SMBH). The most accurate way to treat the simultaneous action of all of these processes is with high-resolution direct $N$-body simulations, which integrate the exact gravitational forces without relying on the orbit-averaged VRR approximation; however, direct $N$-body simulations of the full NSC remain computationally prohibitive – state-of-the-art models of the entire nuclear cluster require months to complete, even on modern GPUs (Panamarev et al., 2019; Panamarev and Kocsis, 2022).
@@ 2 Secular VRR Hamiltonian and Asymptotic Limits
We begin by considering a system of $N$ objects moving on apsidally precessing Keplerian orbits around a central SMBH, with fixed semi-major axes $a_{i}$ and eccentricities $e_{i}$. The central mass is denoted $M$. We assume that $\sum_{i}m_{i}\ll M$, so the gravitational forces from the stars on each other can be treated as small perturbations on otherwise Keplerian orbits. General relativistic corrections and external torques (e.g. from a massive perturber or frame-dragging) are neglected or assumed sub-dominant on the time-scales of interest, unless stated otherwise. Within the VRR framework, over time-scales long compared to an orbital period but short compared to changes in the scalar angular momenta $L_{i}$, the orbits can, therefore, be treated as apsidally precessing annuli whose mutual gravitational interaction is suitably averaged (see e.g. Kocsis and Tremaine, 2015). To build an intuitive understanding of these interactions, we begin by analysing the mutual torques between two circular rings; we then describe the Hamiltonian of the system, which provides a compact and general framework for describing their long-term secular evolution.
@@ 2.1 Torques between rings
where $\eta_{1}=m_{1}/2\pi R_{1}$ and $\eta_{2}=m_{2}/2\pi R_{2}$ are the linear mass densities of ring 1 and ring 2, respectively. Here, $m_{1}$ and $m_{2}$ are the total masses of rings 1 and 2, respectively. The integrals are taken over $\phi_{1}$ and $\phi_{2}$, which represent the azimuthal angles in the coordinate systems corresponding to ring 1 and ring 2. This equation reveals that co-planar rings ($I=0$) exert no gravitational torque on each other because the configuration is symmetric. Similarly, for polar orbits, when $I=\pi/2$, the integral cancels out due to equal and opposite contributions from points on opposite sides of ring 2; the system’s symmetry also leads to $\left(\tau_{12}\right)_{z}=0$. Thus, if we now take the rings to represent a star and an IMBH, with $m_{2}\gg m_{1}$, the $x$-component of the torque induces nodal precession of the star’s angular momentum vector around that of the IMBH.
While Eq. (1) offers an intuitive understanding of the gravitational interactions between two rings, extending this analysis to the $N$-ring system as a whole, and generalising for elliptical orbits, are more conveniently achieved through the Hamiltonian formalism. The gravitational interaction-energy of the ring $i$ with all other rings (Kocsis and Tremaine, 2015) can be written as:
@@
$\mathcal{J}_{ij\ell}=\frac{Gm_{i}m_{j}}{a_{\rm out}}[P_{\ell}(0)]^{2}\,s_{ij\ell}\,\alpha_{ij}^{\ell}\,$, (3)
where $\alpha_{ij}\equiv a_{\mathrm{in}}/a_{\mathrm{out}}$, $a$ is the semi-major axis with indices $\mathrm{out}$}’ and $\mathrm{in}$}’ denoting the larger and smaller semi-major axes among $i$ and $j$, respectively; the coefficient $s_{ij\ell}$ is:
@@ 2.2 Equation of motion and asymptotic limits
From the Hamiltonian (2), one can get the equation of motion (Kocsis and Tremaine, 2015):
@@
$\Omega_{i,j}$ $=2\pi\omega_{i}\frac{m_{j}}{M}\left(\frac{r_{i}}{r_{\mathrm{out}}}\right)\left(\frac{r_{\mathrm{in}}}{r_{\mathrm{out}}}\right)^{2}[P_{2}(0)]^{2}P^{\prime}_{2}(\cos{\theta})$
@@
$M$ is the mass of the SMBH, $\Omega_{i,j}\neq\Omega_{j,i}$, but $\Omega_{i,j}\times L_{i}=\Omega_{j,i}\times L_{j}$, and $\omega_{i}=(GM/a_{i}^{3})^{1/2}$ is the orbital frequency around the SMBH and $\theta$ is the angle between the angular-momentum vectors given by $\cos\theta=\hat{\bm{L}}_{i}\cdot\hat{\bm{L}}_{j}$. This expression is a good approximation when $r_{\rm in}/r_{\rm out}\lesssim 0.3$, as we show in appendix C. One can use an asymptotic expansion (as $\ell\to\infty$) to derive an accurate approximation for arbitrary $r_{\rm in}/r_{\rm out}$; we do so explicitly in Appendix C.
@@
On the other hand, for orbits with close separations which include overlapping or marginally overlapping orbits, large $\ell$ harmonics dominate, because the coefficient $\mathcal{J}_{ij\ell}$ decays slowly with $\ell$, and the torques can be approximated by asymptotic expressions for $\mathcal{J}_{ij\ell}$ at $\ell\to\infty$. As we show in Appendices A and B (cf. Kocsis and Tremaine, 2015) an asymptotic expression valid for both overlapping ($z=1$) and non-overlapping orbits ($z<1$) is
@@
$=\left\{\begin{array}[]{ll}\dfrac{m_{j}a_{i}}{\pi^{2}Mr_{\rm p,out}}\dfrac{[(1+e_{\rm in})(1-e_{\rm out})]^{3/2}}{(e_{i}e_{j})^{1/2}(1-e_{i}^{2})^{1/2}}&\hskip-2.0pt\text{if non-overlapping}\\[8.61108pt] \dfrac{4}{\pi^{3}}\dfrac{m_{j}}{M}\dfrac{I_{2}}{a_{j}(1-e_{i}^{2})^{1/2}}&\hskip-2.0pt\text{if overlapping}\end{array}\right$. (14)
Here, $I_{2}\equiv I^{(2)}(r_{\rm p,in},r_{\rm p,out},r_{\rm a,in},r_{\rm a,out})$ is an analytic function expressed with elliptic integrals in a closed form, given by Eq. (B68) of Kocsis and Tremaine (2015) which simplifies for the case of a circular ring embedded within an elliptical one, as we show in Appendix C.2. In Appendix C.1 we show how to construct equations of motion valid for arbitrary values of radial separation of the orbits.
In § 3 below, we will use these asymptotic expansions, together with exact low-$\ell$ terms, to derive and compare analytical torque-estimates in systems influenced by an additional massive perturber.
@@ 3 Relative torques
[…]
@@
[…]
@@
[…]
@@
$\times|\bm{L}_{i}+\bm{L}_{j}|\sin\theta_{\bullet},$ (18)
where $\kappa\leq 1$ is a geometric factor of order unity which accounts for the projection of the difference vector $\Delta\hat{\bm{L}}$ onto the gradient of the precession frequency. Since the precession frequency $\Omega_{\bullet}$ depends on the inclination $\theta$ relative to $\bm{L}_{\bullet}$, the differential torque is maximized when $\Delta\hat{\bm{L}}$ is aligned with the direction of changing $\theta$ (i.e., along a meridian of the sphere defined by the pole $\bm{L}_{\bullet}$; see Fig. 1). Conversely, the effect vanishes if $\Delta\hat{\bm{L}}$ is azimuthal (along a latitude). $\kappa$ captures the time-averaged orientation of the precessing pair $(\bm{L}_{i},\bm{L}_{j})$ with respect to this frame; its effective value can be calibrated using N-ring simulations. The geometry of the system gives constrains on the maximum relative inclination of the pair – when the system remains coupled, the relative angle is constant; but when it is decoupled, the relative angle varies between $\cos(\theta_{b})=\cos(\max[\theta_{1}\pm\theta_{2}])$ where “max” stands for the configuration where the initial polar angles $\theta_{1,2}$ are aligned according to a meridian as in Figure 1(a).
@@
We now restrict our attention to a special – but analytically tractable – configuration, in which the two bodies share the same semi-major axis and eccentricity, viz. $a_{i}=a_{j}=a$ and $e_{i}=e_{j}=e$. In such a synchronous pair, disruption can arise only from changes in their mutual inclination. The disruption criterion therefore reduces to
$\Omega_{i,j}\!\left(\theta_{b},1\right)m_{i}\leq\;\frac{1}{2}\kappa\,\left|\sin\theta_{\bullet}\,\frac{\partial\Omega_{\bullet}\!\left(\theta_{\bullet},a/a_{\bullet}\right)}{\partial\theta}\right|(m_{i}+m_{j}),$ (19)
@@
When the differential torque due to the IMBH significantly exceeds the mutual torque of the pair, the disruption time-scale can be estimated as
$t_{\rm dis}(a,\theta_{\bullet})\simeq\frac{\pi}{|\Omega_{i,\bullet}-\Omega_{j,\bullet}|}\;\approx\;\frac{\pi}{\theta_{b}\,\bigl|\partial\Omega_{\bullet}\!\left(\theta_{\bullet},a/a_{\bullet}\right)/\partial\theta\bigr|},$ (20)
which corresponds to the time required for the relative precession phase around the IMBH to drift by $\Delta\phi\simeq\pi$, i.e. for the pair to reach its maximum angular separation on the sphere.